Please help me solve this and can you show me how you did it? To rectangles are similar. The ratio of their corresponding sides is 1:4. Find the ratio of there perimeters and their areas
Found 3 solutions by richwmiller, drk, solver91311:
Answer by richwmiller(17219)

If their ratio is 1:4 that means every thing on one is 4 times the size of the other
so one will have the p=2L+2W and A=L*W
the other will be p=2(4L) +2*(4W) and A=4L*4W
p=8L +8W so the ratio of perimeter will be 2:8=1:4
A=16*L*W will be 1:16 or 1:4^2

Answer by drk(1908)

We start with ratio of corresponding sides as 1:4.
Sides, lengths, widths, and perimeters are all called 1 dimensional.
area, surface area, lateral area are all called 2 dimensional.
Volumes are called 3 dimensional
--
To go from a 1 dimensional ratio to a 1 dimensional ratio, just use the same fraction.
To go from a 1 dimensional ratio to a 2 dimensional ratio, square the 1 dimensional fraction.
To go from a 1 dimensional ratio to a 3 dimensional ratio, cube the 1 dimensional fraction
--
we want the ratio of perimeter, so 1:4 = 1:4.
We want the area ratio, so 1:4 - -> (1:4)^2 = 1:16.

Answer by solver91311(24713)



If the ratio of the corresponding sides is 1:4, then if the measure of the length of the smaller rectangle is , then the measure of the length of the larger rectangle is . Likewise, the measures of the widths must be and .

The perimeter of the small rectangle is:



and the perimeter of the large rectangle is:



From the last relationship, factor out a 4:



Note that the quantity inside the parentheses is identical to the RHS of the first equation, so:



Hence, the ratio of is

The area of the smaller rectangle is



The area of the larger rectangle is



Ratio is


John